# Optimus: Accelerating Large-Scale Multi-Modal LLM Training by Bubble Exploitation

### v1

@@ Abstract
Multimodal large language models (MLLMs) have extended the success of large language models (LLMs) to multiple data types, such as image, text and audio, achieving significant performance in various domains, including multimodal translation, visual question answering and content generation. Nonetheless, existing systems are inefficient to train MLLMs due to substantial GPU bubbles caused by the heterogeneous modality models and complex data dependencies in 3D parallelism. This paper proposes Optimus, a distributed MLLM training system that reduces end-to-end MLLM training time. Optimus is based on our principled analysis that scheduling the encoder computation within the LLM bubbles can reduce bubbles in MLLM training. To make scheduling encoder computation possible for all GPUs, Optimus searches the separate parallel plans for encoder and LLM, and adopts a bubble scheduling algorithm to enable exploiting LLM bubbles without breaking the original data dependencies in the MLLM model architecture. We further decompose encoder layer computation into a series of kernels, and analyze the common bubble pattern of 3D parallelism to carefully optimize the sub-millisecond bubble scheduling, minimizing the overall training time. Our experiments in a production cluster show that Optimus accelerates MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B model over 3072 GPUs compared to baselines.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) continue the hot of Large Language Models (LLMs) and further extend LLM’s capability to understand and generate content from multiple modalities (e.g., text, images, and audio). MLLMs, such as GPT-4V [22], Google Gemini [29], Grok-1.5 Vision [33] and LLava [19], have achieved remarkable progress in various domains, such as visual question answering [2, 20], multimodal translation [28, 34], and content generation and understanding [29, 22, 39]. Notably, the computational demands of MLLMs are substantial, emphasizing the urgent need to enhance training performance to fully leverage their capabilities.
MLLMs typically involve the integration of multiple encoders, each tailored to process specific modalities, combined with a giant language model component. The multimodal data is passed to respective encoders, and the output is combined to serve as the input of the language model.
@@
The multimodal encoders and the language model vary greatly in functionalities, architectures, and data input sizes, leading to different resource demands. However, existing distributed training systems are mainly designed for sequential unimodal (e.g., MegaScale [14], Megatron-LM [21], Chimera [17]), and fall short in MLLMs training with over 40% idle GPU cycles when we train a large MLLM (several hundred of billions of parameters) using Megatron-LM and more than 3,000 GPUs. After analyzing typical MLLM training tasks, we made two key observations. (1) The communication of 3D parallelism is extensive and frequent, leading to long GPU idle time. (2) The pipeline stages of MLLM are imbalanced and the data dependency between adjacent pipeline stages results in long data waiting time. Existing solutions can be classified into two categories: (1) optimizing LLM, e.g., Megatron-LM and Zero-bubble pipeline[24]; (2) optimizing multimodal encoders, e.g., DistMM[13]. Nonetheless, none of the existing works consider LLM and encoders together and we will show in Section 2.2 that around 48% GPU cycles are wasted in our internal large-scale MLLM training task.
@@
In this paper, we propose Optimus, a distributed MLLM training system that enables the scheduling of encoder computation within LLM bubbles to achieve performant 3D parallelism. However, it is difficult to schedule encoder computation within LLM bubbles based on existing training frameworks because of three main reasons.
@@
First, existing training frameworks, e.g., Megatron-LM [21], MegaScale [14], and zero-bubble pipeline [24], apply unified parallel strategies to MLLM models, distributing encoder and LLM layers across different GPUs. As a result, most GPUs contain only LLM model states, unable to perform encoder computation during LLM bubbles. In contrast, we use separate parallel plans for encoders and LLM to colocate encoder and LLM model states on each GPU. We enumerate potential 3D parallelism plans for the encoder and prune plans that violate the GPU memory constraint.
@@
Second, the presence of complex data dependencies within MLLM imposes constraints on the scheduling of encoder computation within LLM bubbles. There are dependencies related to synchronous training iterations and internal dependencies within the encoder (see Section 2.3). The most intricate of these is the encoder-LLM microbatch-level data dependency, which necessitates that the encoder completes its forward pass before the LLM begins its forward pass for each microbatch and that the encoder begins its backward pass only after the LLM has completed its backward pass for each microbatch. To manage these dependencies, we employ a two-stage dependency management approach: local scheduling to address the first two types of dependencies and global ordering to handle the encoder-LLM microbatch-level dependencies.
@@
Third, the LLM bubble duration varies from sub-milliseconds to hundreds of milliseconds, making bubble reduction a hurdle to overcome. Existing frameworks [21, 24, 17] schedule in the unit of layers, and the sub-millisecond bubble is too short to complete even a single encoder layer forward or backward. Hence, we decompose encoder layer computation into a series of kernels to utilize the sub-millisecond bubbles. Further, we analyze the common patterns of LLM bubbles, and optimize the bubble schedule by scheduling encoder kernel computation to bubbles interleaved with LLM computation to minimize the overall training time.
@@
We have implemented Optimus based on Megatron-LM, including the above design points. We conduct extensive experiments using multiple representative MLLM models. The results are promising - Optimus outperforms state-of-the-art baselines by 20.3% on average and Optimus also scales well with the size of models and GPUs. Our experiments in a production cluster show that Optimus accelerates MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B model over 3072 GPUs compared to baselines.
@@ 2.1 Multimodal LLM Characteristics
Multimodal LLMs are increasingly important. These models inherit the foundational principles of LLMs, integrating advanced natural language processing techniques while expanding their scope to encompass diverse data modalities. GPT-4[22] represents a prominent example of a multimodal model that extends the capabilities and success of its predecessors to encompass multimodal understanding and generation, demonstrating human-level performance in various benchmark tests with inputs of both images and text.
Multimodal large language model (MLLM) comprises three key parts: one or multiple modality encoders, input projectors, and a large language model backbone[36]. The Modality Encoders are designed to encode inputs from non-textual modalities into respective features, while the input projector aligns features from these modalities with the text feature space. Ultimately, the LLM backbone utilizes aligned features from various modalities and textual features as its input. Figure 1 illustrates the architecture of the MLLM. We exclude the input projector from our discussion due to its relatively minor computational demand compared to the encoder and the LLM (refer to Llava [19]). Additionally, we treat the input projector as the final layer of the modality encoder in our analysis.
Figure 1: Multimodal model architecture
@@ 2.2 Bubbles in MLLM Training
Existing LLM pipeline optimizations are not model-agnostic, and fall short in MLLM training tasks. In our internal large-scale MLLM training tasks with ViT encoder and GPT backbone (over 100B parameters), we train Megatron-LM with more than 3,000 NVIDIA GPUs and observe more than 48% GPU cycle idleness when applying multiple SOTA techniques, including MegaScale [14], Zero Bubble Pipeline [24], fine-grained communication-computation overlapping [32]. We analyze the profiled timeline to identify and investigate the occurrences of GPU idleness (i.e., bubbles). Table 1 shows the total time and percentage of average training step time (5.12s) occupied by different types of bubbles.
@@
Table 1: Total time and percentage of average training step time (5.12s) occupied by different types of bubbles
@@
(1) Communication in Data Parallelism (DP). Data parallelism requires communication to aggregate gradients, leading to GPU idle time during the communication. Specifically, MegaScale [14] and Megatron-LM [26] use the distributed optimizer (similar to $P_{os+g}$ in ZeRO [25]) to save memory for large model training, which performs two collective communications (all-gather and reduce-scatter). At the start of each training step, an all-gather operation gathers updated parameters from all data parallel (DP) ranks, resulting in a DP all-gather bubble (occupying 3.3% of the training time). At the end of the training step, reduce-scatter is performed to aggregate gradients, leading to a DP reduce-scatter bubble (occupying 8.9% of the training time). It should be noted that overlapping optimization in data parallelism proposed in Megascale [14] have already been applied and above DP communications are required for the first model chunk which can not be hidden because of the nature of synchronous training  [14]. Figure 2 illustrates DP bubbles that occur due to all-gather and reduce-scatter operations at the start and conclusion of each training step.
@@
Figure 2: Timeline illustration of MLLM training showing different categories of bubbles (simplified based on the Megatron-LM 1F1B schedule [21])
@@
(2) Dependency in Pipeline Parallelism (PP). Despite applying pipeline send-receive overlap optimization from Megascale [14], pipeline bubbles still occur due to the inherent data dependencies between stages during the forward and backward passes. It should be noted that Zero Bubble Pipeline cannot eliminate pipeline bubbles in MLLM training, owing to the required changes in the optimizer [24] (refer to discussions in §7). Figure 2 illustrates the MLLM training pipeline schedule, which consists of three phases: warm-up (forward only), steady (one forward and one backward), and cool-down (backward only). Throughout pipeline training, three types of bubbles arise:
@@
  1. PP warm-up bubbles occur at all stages except the initial one due to the forward dependency of the first forward pass, averaging 5.0% of the training time.
@@
  2. PP cool-down bubbles occur at all stages except the initial one due to the backward dependency of the final backward pass, averaging 9.2% of the training time.
@@
  3. Other PP bubbles manifest at all stages except the last one due to dependencies of other forward and backward passes, occupying 8.7% of training time. For instance, PP bubbles emerge immediately after the PP warm-up phase due to the backward dependency of the initial backward pass. Additionally, in cases of imbalanced pipeline stages caused by MLLM’s heterogeneous model, there are additional pipeline bubbles not depicted in Figure 2.
@@
(3) Communications in Tensor Parallelism (TP). Tensor parallelism entails partitioning individual model layers across multiple GPUs, necessitating communication during forward and backward passes to synchronize between GPUs. In Megatron-LM, each forward or backward pass of a transformer layer involves two all-gather and two reduce-scatter kernels [15]. Figure 3 provides a detailed view of CUDA computation and communication kernels during two GPT-175B[6] layer forward passes. In the CUDA communication stream, green kernels represent all-gather communications, while blue kernels denote reduce-scatter communications. The compute stream idles during these communications. Typically, these TP bubbles last for sub-millisecond durations, averaging around 300 $\mu$s. However, during MLLM training, there are thousands of TP bubbles, totaling 11.2% of the training time.
Figure 3: Zoom-in view of TP bubbles during two LLM layers forward
@@ 2.3 Challenges
To minimize bubbles in MLLM training, we aim to leverage the distinct dual-component structure of MLLM, which includes encoders and the LLM backbone. We have noted two key observations. Firstly, the majority of bubbles during MLLM training tend to occur during the forward and backward passes of the LLM backbone, with around 90% of these bubbles arising from LLM communication, as indicated in Table 1. Secondly, the encoders require fewer computational operations (FLOPs) than the LLM backbone due to their smaller number of parameters[5, 8, 19, 18, 11].
In response, we propose to schedule encoder computation in LLM bubbles (occurring during communication in LLM) to reduce bubbles throughout the MLLM training process.
We identify three main challenges of scheduling encoder computation to LLM bubbles.
Challenge 1: Only a few GPUs have both encoder and LLM model states. Current training systems [38, 21] use pipeline parallelism to parallelize the MLLM as a single pipeline. Due to the dependency between the encoder and LLM, encoder layers are assigned to earlier pipeline stages, while LLM layers are assigned to later pipeline stages. Consequently, only one pipeline stage typically contains both encoder and LLM layers. To illustrate, Figure 4 demonstrates the application of 3D parallelism (DP=1, PP=4, TP=2) to parallelize MLLM across 8 GPUs, where only 2 GPUs in pipeline stage 1 possess both encoder and LLM model states. The remaining 6 GPUs are incapable of executing encoder computations during LLM bubbles because they lack encoder model states.
Figure 4: Only GPUs in pipeline stage 1 have both encoder and LLM model states
Challenge 2: Complex Dependencies in MLLM Training. The intricate dependencies inherent in MLLM training pose significant challenges when scheduling encoder computation within LLM bubbles. Firstly, in synchronous training, the utilization of LLM bubbles is restricted to executing the required encoder computation solely within the current training iteration (iteration dependency). Secondly, the dependency within the encoder pipeline requires scheduling the forward computation of the current encoder pipeline stage $i$ after the completion of the previous encoder stage, and scheduling the backward computation after the subsequent encoder stage concludes. Lastly, the encoder-LLM dependency entails a microbatch-level dependency, where the encoder must complete the forward pass of microbatch $i$ before the LLM pipeline initiates the forward pass of microbatch $i$, and similarly, the encoder can commence the backward pass of microbatch $i$ after the LLM pipeline completes the backward pass of microbatch $i$.

### v2

@@ Abstract
Multimodal large language models (MLLMs) have extended the success of large language models (LLMs) to multiple data types, such as image, text and audio, achieving significant performance in various domains, including multimodal translation, visual question answering and content generation. Nonetheless, existing systems are inefficient to train MLLMs due to substantial GPU bubbles caused by the heterogeneous modality models and complex data dependencies in 3D parallelism. This paper proposes Optimus, a distributed MLLM training system that reduces end-to-end MLLM training time. Optimus is based on our principled analysis that scheduling the encoder computation within the LLM bubbles can reduce bubbles in MLLM training. To enable scheduling encoder computation for all GPUs, Optimus searches for separate parallel plans for the encoder and LLM, and adopts a bubble scheduling algorithm to exploit LLM bubbles without breaking the original data dependencies in the MLLM model architecture. We further decompose the encoder layer computation into a series of kernels and analyze the common bubble pattern of 3D parallelism to carefully optimize the sub-millisecond bubble scheduling, minimizing the overall training time. Our experiments in a production cluster show that Optimus accelerates MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B model over 3072 GPUs compared to baselines.
**footnotetext: Work done during internship at Bytedance.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) build upon the advancements of Large Language Models (LLMs) by extending their capabilities to comprehend and generate content across multiple modalities, such as text, images, and audio. Notable MLLMs, including GPT-4V [21], Google Gemini [28], Grok-1.5 Vision [33] and LLava [18], have achieved remarkable progress in domains like visual question answering [2, 19], multimodal translation [27, 34], and content generation and understanding [28, 21, 40]. The substantial computational demands of MLLMs underscore the urgent need to enhance training performance to fully harness their capabilities.
Multimodal large language models (MLLMs) typically integrate multiple encoders, each specialized for processing a specific modality, alongside a substantial language model component. Multimodal data are input into their respective encoders, and the resulting outputs are combined to form the input for the language model.
@@
The multimodal encoders and the language model exhibit significant differences in functionality, architecture, and input data sizes, leading to varied resource demands. However, existing distributed training systems are mainly designed for sequential unimodal (e.g., MegaScale [13], Megatron-LM [20], Chimera [16]), and fall short in training MLLMs. For example, when training a large MLLM containing several hundred billion parameters using Megatron-LM on over 3,000 GPUs, more than 40% of GPU cycles remain idle. Upon analyzing typical MLLM training tasks, we identified two critical issues: (1) Communication overhead in 3D parallelism is extensive and frequent, resulting in significant GPU idle times; (2) The pipeline stages of MLLM are imbalanced, and the data dependency between adjacent pipeline stages results in considerable data waiting time. Existing solutions can be categorized into two groups: (1) optimizing LLM, e.g.Megatron-LM and Zero-bubble pipeline[23]; (2) optimizing multimodal encoders, e.g.DistMM[12]. Nonetheless, none of the existing works consider LLM and encoders simultaneously. As demonstrated in Section 2.2, around 48% of GPU cycles are wasted in our internal large-scale MLLM training task.
@@
In this paper, we introduce Optimus, a distributed MLLM training system that enables scheduling encoder computations within idle periods — referred to as "LLM bubbles" — to achieve efficient 3D parallelism. Scheduling encoder computations within these LLM bubbles using existing training frameworks is challenging for three reasons.
@@
First, existing training frameworks, e.g., Megatron-LM [20], MegaScale [13], and zero-bubble pipeline [23], employ unified parallel strategies to MLLM models, distributing encoder and LLM layers across different GPUs. As a result, most GPUs contain only LLM model states and are incapable of performing encoder computations during LLM bubbles. In contrast, our approach utilizes separate parallelization plans for encoders and LLMs to colocate both encoder and LLM model states on each GPU. We systematically enumerate potential 3D parallelism plans for the encoder and eliminate those that violate GPU memory constraints.
@@
Second, complex data dependencies inherent in MLLM constrain the scheduling of encoder computations within LLM bubbles. These dependencies include those related to synchronous training iterations and internal dependencies within the encoder itself, as discussed in Section 2.3. The most intricate dependency is the microbatch-level data dependency between the encoder and LLM, which requires that the encoder completes its forward pass before the LLM begins its forward pass for each microbatch, and that the encoder’s backward pass commences only after the LLM has completed its backward pass for each microbatch. To manage these dependencies, we employ a two-stage dependency management strategy: local scheduling to address the first two types of dependencies and global ordering to handle the encoder-LLM microbatch-level dependencies.
@@
Third, the LLM bubble duration varies significantly, ranging from sub-millisecond intervals to several hundred milliseconds, posing a considerable challenge for bubble reduction. Existing frameworks [20, 23, 16] schedule computations at the layer level, and sub-millisecond bubbles are too brief to complete even a single encoder layer forward or backward. To address this, we decompose encoder layer computations into sequences of kernels, enabling the effective utilization of these brief bubbles. Furthermore, we analyze common patterns in LLM bubbles and optimize the scheduling by interleaving encoder kernel computations with LLM computations, thereby minimizing the overall training time.
@@
Optimus was implemented based on Megatron-LM, incorporating the aforementioned design principles. Comprehensive evaluations were conducted using several representative MLLM models. The results are promising - Optimus outperforms state-of-the-art baselines by 20.3% on average and Optimus also scales well with the size of models and GPUs. Our experiments in a production cluster show that Optimus accelerates MLLM training by 20.5%-21.3% with ViT-22B and GPT-175B models over 3072 GPUs compared to baselines.
@@ 2.1 Multimodal LLM Characteristics
Multimodal LLMs are increasingly important. These models build upon the foundational principles of LLMs by integrating advanced natural language processing methodologies while extending their scope to encompass diverse data modalities. GPT-4[21] exemplifies such an advancement, enhancing the capabilities of its predecessors to include multimodal understanding and generation. It demonstrates human-level performance across various benchmark tests involving both image and text inputs.
A multimodal large language model (MLLM) typically comprises three primary components: one or more modality encoders, input projectors, and a large language model backbone [36]. The Modality Encoders process inputs from non-textual modalities into their respective feature representations, while the input projector aligns these features with the text feature space. Subsequently, the LLM backbone utilizes the aligned multimodal and textual features as input. Figure 1 illustrates the architecture of the MLLM. The input projector is excluded from further discussion due to its relatively minimal computational demands compared to the encoder and the LLM backbone, as detailed in Llava [18]. For this analysis, the input projector is considered the final layer of the modality encoder.
Figure 1: Multimodal model architecture.
@@ 2.2 Bubbles in MLLM Training
Existing LLM pipeline optimizations lack model-agnostic properties and are inadequate for MLLM training tasks. In our internal large-scale MLLM training tasks with ViT encoder and GPT backbone (over 100B parameters), we employed Megatron-LM across more than 3,000 NVIDIA GPUs and observed over 48% GPU cycle idleness despite utilizing multiple SOTA techniques, including MegaScale [13], Zero Bubble Pipeline [23], fine-grained communication-computation overlapping [32]. We analyzed the profiled timeline to identify and investigate the occurrences of GPU idleness (i.e., bubbles). Table 1 shows the total duration and percentage of average training step time (5.12s) occupied by different types of bubbles.
@@
Table 1: Total time and percentage of average training step time (5.12s) occupied by different types of bubbles.
@@
(1) Communication in Data Parallelism (DP). Data parallelism requires communication to aggregate gradients, leading to GPU idle time during these communications. Specifically, MegaScale [13] and Megatron-LM [25] employ a distributed optimizer, similar to $P_{os+g}$ in ZeRO [24], to save memory for large model training, which performs two collective communications: all-gather and reduce-scatter. At the start of each training step, an all-gather operation collects updated parameters from all DP ranks, resulting in a DP all-gather bubble that occupies 3.3% of the training time. At the end of the training step, reduce-scatter is performed to aggregate gradients, leading to a DP reduce-scatter bubble that consumes 8.9% of the training time [^1] The reduce-scatter bubble is larger than the all-gather bubble because it involves higher communication volume (FP32 gradients vs. BF16 parameters) and suffers from synchronization delays due to straggling GPUs at the end of the training step.. It should be noted that overlapping optimization in data parallelism proposed in Megascale [13] has already been applied; however, the aforementioned DP communications for the first model chunk can not be hidden due to the nature of synchronous training [13]. Figure 2 illustrates DP bubbles arising from all-gather and reduce-scatter operations at the start and conclusion of each training step.
@@
Figure 2: Timeline illustration of MLLM training showing different categories of bubbles (simplified based on the Megatron-LM 1F1B schedule [20]).
@@
(2) Dependency in Pipeline Parallelism (PP). Despite applying pipeline send-receive overlap optimization from Megascale [13], pipeline bubbles persist due to the inherent data dependencies between stages during the forward and backward passes. Importantly, the Zero Bubble Pipeline method cannot eliminate pipeline bubbles in MLLM training, owing to the required changes in the optimizer [23] (see discussions in §7). Figure 2 illustrates the MLLM training pipeline schedule, comprising three phases: warm-up (forward only), steady (one forward and one backward), and cool-down (backward only). Throughout pipeline training, three types of bubbles arise:
@@
  1. PP warm-up bubbles occur at all stages except the first one due to the forward dependency of the first forward pass, averaging 5.0% of the training time.
@@
  2. PP cool-down bubbles occur at all stages except the first one due to the backward dependency of the final backward pass, averaging 9.2% of the training time.
@@
  3. Other PP bubbles manifest in all stages except the last one due to dependencies of other forward and backward passes, occupying 8.7% of training time. For instance, PP bubbles emerge immediately after the PP warm-up phase due to the backward dependency of the initial backward pass. Additionally, in cases where pipeline stages are imbalanced due to the heterogeneity of MLLM models, extra pipeline bubbles not depicted in Figure 2 occur.
@@
(3) Communications in Tensor Parallelism (TP). Tensor parallelism involves partitioning individual layers across multiple GPUs, necessitating communication during forward and backward passes to synchronize between GPUs. In Megatron-LM, each forward or backward pass of a transformer layer requires two all-gather and two reduce-scatter kernels [14]. Figure 3 provides a detailed view of CUDA computation and communication kernels during two GPT-175B [5] layer forward passes. Green kernels represent all-gather communications in the CUDA communication stream, while blue kernels denote reduce-scatter communications. The compute stream idles during these communications. Typically, these TP bubbles last for sub-millisecond durations, averaging around 300 $\mu$s. However, during MLLM training, there are thousands of TP bubbles, collectively accounting for 11.2% of the training time.
Figure 3: Zoom-in view of TP bubbles during two LLM layers forward.
@@ 2.3 Challenges
To minimize bubbles in MLLM training, we aim to exploit the distinct dual-component structure of MLLM, which consists of encoders and the LLM backbone. We make two key observations. First, the majority of bubbles during MLLM training occur during the forward and backward passes of the LLM backbone, with around 90% arising from LLM communication based on production data. Second, the encoders require fewer computational operations (FLOPs) than the LLM backbone due to their smaller parameter size[4, 7, 18, 17, 10].
Based on these insights, we propose scheduling encoder computations in LLM bubbles which occur during LLM communication to minimize bubbles throughout the MLLM training process. We identify three primary challenges of scheduling encoder computation to LLM bubbles.
Challenge 1: Only a limited number of GPUs have both encoder and LLM model states. Current training systems [39, 20] employ pipeline parallelism to distribute the MLLM as a single pipeline across multiple GPUs. Due to the dependency between the encoder and LLM, encoder layers are assigned to initial pipeline stages, while LLM layers are assigned to later pipeline stages. Consequently, only one pipeline stage typically contains both the encoder and LLM layers. To illustrate, Figure 4 demonstrates the use of 3D parallelism (DP=1, PP=4, TP=2) to parallelize MLLM across 8 GPUs, where only 2 GPUs in pipeline stage 1 hold both the encoder and LLM model states. The remaining 6 GPUs cannot execute encoder computations during LLM bubbles because they lack encoder model states.
Figure 4: Only GPUs in pipeline stage 1 have both encoder and LLM model states.
Challenge 2: Complex Dependencies in MLLM Training. The complex dependencies inherent in MLLM training pose significant challenges when scheduling encoder computation within LLM bubbles. First, in synchronous training, the utilization of LLM bubbles is restricted to performing encoder computations solely within the current training iteration (iteration dependency). Second, dependencies within the encoder pipeline require that the forward computation of the current encoder pipeline stage $i$ is scheduled only after the completion of the previous encoder stages, and the backward computation can be scheduled only after the subsequent encoder stage concludes. Lastly, the encoder-LLM dependency imposes microbatch-level constraints: the encoder must complete the forward pass of microbatch $i$ before the LLM pipeline can initiate the forward pass of the same microbatch. Similarly, the encoder can begin the backward pass of microbatch $i$ only after the LLM pipeline has completed its backward pass of microbatch $i$.